# The compact object of HESS J1731-347 and its implication on neutron star matter

### v1

@@ V Conclusion
While we were conducting this study, a new simultaneous mass-radius measurement of the nearest and the brightest milliseond pulsar PSR J0437-4715 was published by the NICER collaboration Choudhury et al. (2024). The new data matches well with our results at the $90\%$ CI. Hence, we have not incorporated it into our analysis. However, we plan to conduct a thorough study to quantify the effect of the new data.
@@ Acknowledgements
This work has been partially supported by the Fonds de la Recherche Scientifique-FNRS, Belgium, under grant No. 4.4501.19. PC is supported by European Union’s HORIZON MSCA-2022-PF-01-01 Programme under Grant Agreement No. 101109652, project ProMatEx-NS. BB acknowledges the support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2121 “Quantum Universe” – 390833306.

### v2

@@ V Conclusion
While we were conducting this study, a new simultaneous mass-radius measurement of the nearest and the brightest millisecond pulsar PSR J0437-4715 was published by the NICER collaboration Choudhury et al. (2024). The new data matches well with our results at the $90\%$ CI. Hence, we have not incorporated it into our analysis. However, we plan to conduct a thorough study to quantify the effect of the new data.
@@ Acknowledgements
This work has been partially supported by the Fonds de la Recherche Scientifique-FNRS, Belgium, under grant No. 4.4501.19. PC is supported by European Union’s HORIZON MSCA-2022-PF-01-01 Programme under Grant Agreement No. 101109652, project ProMatEx-NS. BB acknowledges the support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2121 “Quantum Universe” – 390833306 and the Alexander von Humboldt Foundation through a Humboldt Research Fellowship for Postdoctoral Researchers.